OpenStudy (anonymous):

2. Arianna kicks a soccer ball off the ground and into the air with an initial velocity of 42 feet per second. Assume the starting height of the ball is 0 feet. Approximately what maximum height does the soccer ball reach? (1 point) 1.3 ft 2.6 ft 26.0 ft 27.6 ft 3. Arianna kicks a soccer ball off the ground and into the air with an initial velocity of 42 feet per second. Assume the starting height of the ball is 0 feet. Approximately, how long does it take until the soccer ball hits the ground again? (1 point) 1.3 secs 2.6 secs 16.0 secs 27.6 secs

OpenStudy (anonymous):

well for mechanics there are 3 big equations to be used for motions: \[v = v_{0} + a \times t\] \[x = v_{0} \times t + \frac{ a \times t^{2}}{ 2}\] \[v^{2} = v_{0}^{2} + 2 \times a \times x\]

OpenStudy (anonymous):

when it comes to vertical motion g - gravitational acceleration replaces a, and h - height replaces x ( distance)

OpenStudy (anonymous):

i dont get any of that :(

OpenStudy (anonymous):

I'm really bad at math... i've been struggling lately

OpenStudy (anonymous):

but here you have to use conservation of energy: \[E_{inital} = E_{final} \] initially the ball has only kinetic energy because the initial height is 0 and in the end the ball has stopped moving because it has reached the maximum height so it has only potential energy: \[E_{k} = E_{p} => \frac{ m \times v^{2} }{ 2 } = m \times g \times h\] simplify by m and get:\[h = \frac{ v^{2} }{ 2 \times g } \]

OpenStudy (anonymous):

so im confused.... idk what the answer would be.... :l